OpenStudy (anonymous):

Find the domain of f(x) = ln(7 - x^2)/2 So I solve for this inequality, because the argument of the natural log must be strictly greater than 0. 7 - x^2 > 0 7 > x^2 \sqrt{7} > \sqrt{x^2} <=> \sqrt{7} > |x| Which in interval notation, is: ( -\sqrt{7}, \sqrt{7} ) But this is "wrong" for some reason. Would someone point out why please?

OpenStudy (jamesj):

If you want the domain of \[f(x) = \frac{ \ln(7 - x^2)}{2}\] or even \[f(x) = \ln \left( \frac{7-x^2}{2} \right)\] then what you've done is exactly right. Note in both cases x CAN have negative values. For example, if x = -1, then 7 - x^2 = 6, and we can evaluate either function.

OpenStudy (anonymous):

So this domain on this interval: ( -\sqrt{7}, \sqrt{7} ) , is correct?

OpenStudy (jamesj):

For the problem you've given us, yes. Why do you think it's wrong?

OpenStudy (anonymous):

I entered this on my electronc web assignment and it says that is wrong

OpenStudy (jamesj):

pesky web assignments.